OpenStudy (anonymous):

why is log(-e+e^3) the same as 1+ log(e^2-1)

OpenStudy (zarkon):

I assume the log here is the natural log \[\log(-e+e^3)=\log(e(-1+e^2)=\log(e)+\log(-1+e^2)=1+\log(e^2-1)\]

OpenStudy (anonymous):

Ohhhhhhhhhhhhhh i understand it now THANKS A MILLION @ ZARKON

OpenStudy (zarkon):

No problem :)

OpenStudy (anonymous):

@ zarkon can you also help me solve this \[\int\limits_{1}^{e} (2x+e)/(x^2+ex)\]

OpenStudy (zarkon):

let \[u=x^2+ex\] \[du=(2x+e)dx\]

OpenStudy (anonymous):

yes and what do you get as an answer?

OpenStudy (zarkon):

Looks like it is \[\left.\ln(x^2+ex)\right|_{1}^{e}\] \[\ln(e^2+ee)-\ln(1+e)=\ln(2e^2)-\ln(1+e)\]

OpenStudy (anonymous):

thank you very much!